The bottom of Jim's rectangular bait box is 3 inches longer than it is wide. The diagonal is 15 inches. What is the area of the bottom of the box?
Answer by unlockmath(1688)

Hello,
First you need to find the measurement of of both sides of the rectangle then
we can find the area.
Use the formula A^2 + B^2 = C^2
We know the hypotenuse is 15 inches.
Let x = one side of the rectangle and x + 3 be the other side.
Therefore:
x^2 + (x+3)^2 = 15^2 this works out to
x^2 + x^2 + 6x + 9 = 225 combine like terms
2x^2 + 6x + 9 = 225 subtract 225 from both sides
2x^2 + 6x - 216 = 0 now we can solve x by the quadratic formula
x = 9
Therefore one side is 9 inches and the other side is 12 inches.
Multiply these sides and the answer is:
108 square inches
You're welcome
RJ Toftness
www.math-unlock.com